OpenStudy (anonymous):

a baseball game charges $3 for a ticket and $7 for adults. One game 400 tickets worth $1760 were sold, how many adults and how many children tickets were sold

OpenStudy (johnweldon1993):

Alright, so you know that in all....the sum of children and adult tickets sold is 400 so this can be written as c + a = 400 We know that each child ticket is 3 and each adult is 7...so now we have an equation 3c + 7a = 1760 2 equations...2 unknowns...can you solve from here?

OpenStudy (anonymous):

no, i dont get any of it

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Alright we'll walk through it... c + a = 400 3c + 7a = 1760 Do you see where I got those equations from?

OpenStudy (anonymous):

yea, the tickets sold and how much it came out too.

OpenStudy (johnweldon1993):

Right...so this is called a "system of equations" To solve it we can use a few methods...but the easiest here would be the substitution method.. What this method says to do...is solve one of these equations for a variable...then substitute that value into the other equation to solve for the other... so here...

OpenStudy (johnweldon1993):

c + a = 400 3c + 7a = 1760 Lets solve that top equation for the variable 'c' ... How do we isolate 'c' in that equation? well we subtract 'a' from both sides. c + a = 400 - a - a ------------ c = 400 - a Still with me?

OpenStudy (anonymous):

yea.

OpenStudy (johnweldon1993):

So now lets look at that second equation 3c + 7a = 1760 Well....we just solved and figured out that 'c' = 400 - a....so lets substitute that into this equation...so wherever you see a 'c' ...you replace it with (400 - a) so 3c + 7a = 1760 would become 3(400 - a) + 7a = 1760 what do we do from here?

OpenStudy (anonymous):

400-400 then 400-1760?

OpenStudy (johnweldon1993):

Not quite...so you would distribute the 3 into the parenthesis... so you multiply each term in the parenthesis by 3 3(400 - a) + 7a = 1760 3 times 400 is 1200 3 times -a = -3a so now we have... 1200 - 3a + 7a = 1760 Now what?

OpenStudy (anonymous):

3a + 3a 7a+3a so it would be 1200+10a=1760

OpenStudy (johnweldon1993):

Well remember it is -3a so when we combine the like terms... -3a + 7a is equal to 4a right? because -3 + 7 = 4 So now we have 1200 + 4a = 1760 Next up?

OpenStudy (anonymous):

1200-1200 1200-1760 so 4a=560 4a /4a 4a/560 so a=140

OpenStudy (johnweldon1993):

Right! So we have that 140 Adult tickets were sold. Well, we know that there were 400 total sold...so how many children tickets were sold??

OpenStudy (anonymous):

260

OpenStudy (johnweldon1993):

And there's your answer :) you can check it by plugging both numbers into the equations and solving

OpenStudy (anonymous):

ok, Thank you so much!(:

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

(: